show that a(b^2+c^2)COSA+b(c^2+a^2)COSB+c(a^2+b^2)COSC=3abc?

Grade:12th pass

1 Answers

Avinash Kumar Singh
6 years ago
a^2(b^2+c^2)COSA/aPut value of cosA= b^2 +c^2 - a^2 /2bca^2(b^2+c^2)COSA/a= a^2(b^2+c^2)(b^2 +c^2 - a^2) /2abcAdd all three terms to get 3abc